# DESY 25-127, ZU-TH 59/25, MPP-2025-182, RISC Report Series 25–08 The three-loop single-mass heavy-flavor corrections to the structure functions $F_{2}(x,Q^{2})$ and $g_{1}(x,Q^{2})$

### v1

@@ II The Structure Function $F_{2}(x,Q^{2})$
III The Structure Function $g_{1}(x,Q^{2})$
In the polarized case the calculation is performed in the Larin scheme [54]. Both for the massive three-loop OMEs and the massless three-loop Wilson coefficients, the scheme transformations to the $\overline{\rm MS}$ scheme are not yet known. However, if all contributing quantities, including also the parton distribution functions, are used in this scheme, $g_{1}(x,Q^{2})$, as a scheme independent quantity, can still be assembled. For the process of deep-inelastic scattering, the Larin scheme is a consistent scheme, given the number of contributing $\gamma_{5}$ Dirac matrices. A set of Larin-scheme parton distribution functions has been generated in Ref. [55] and is used in the following illustrations. The massless and massive two-loop corrections for the structure function $g_{1}(x,Q^{2})$ were calculated in Ref. [31].
In Figure 5, we illustrate the scale evolution of the complete structure function $g_{1}^{\rm total}(x,Q^{2})$ from $Q^{2}=25~\rm GeV^{2}$ to $Q^{2}=1000~\rm GeV^{2}$ at three-loop order. With growing values of $Q^{2}$, $xg_{1}^{\rm total}(x,Q^{2})$ grows below $x\sim 0.2$ and depletes at larger values of $x$. Also below $x\sim 10^{-2}$ the order in the scales reverts. The structure functions $g_{1}^{p,d}(x,Q^{2})$ are positive within errors, cf. Ref. [56].
Figure 6: The different heavy-flavor contributions due to charm quarks to the structure function $g_{1}(x,Q^{2})$ due to the Wilson coefficients $\Delta H_{g}^{\rm S},\Delta H_{q}^{\rm PS},\Delta L_{q}^{\rm NS},\Delta L_{g}^{\rm S}$ and $\Delta L_{g}^{\rm PS}$ at $Q^{2}=100~\rm GeV^{2}$ at different cumulative orders in $a_{s}$.
@@ III The Structure Function $g_{1}(x,Q^{2})$
Figure 6 illustrates the contributions of the different heavy-flavor Wilson coefficients to $g_{1}^{c}$ for $Q^{2}=100~\rm GeV^{2}$, also with their contributions at different cumulative order in $a_{s}$. One clearly sees the importance of the higher order corrections. The largest contribution is due to the gluonic Wilson coefficient $\Delta H_{g}^{\rm S}$. As has been discussed in Refs. [29, 31], the first moment of $\Delta H_{g}^{\rm S}(x,Q^{2})$ vanishes,
$\int_{0}^{1}dx\Delta H_{g}^{\rm S}(x,Q^{2})=0.$ (10)
@@
In Figure 8, we show the ratio Eq. (9) of the charm and bottom contributions to the whole structure function $g_{1}(x,Q^{2})$ at NNLO for $Q^{2}=25,100$ and $1000~\rm GeV^{2}$. In the small-$x$ region, the corrections are negative. The largest negative values are about $10~\%,-25~\%$ and $95~\%$ from $Q^{2}=25,100$ to $1000~\rm GeV^{2}$. These large corrections occur in a region in which $xg_{1}(x,Q^{2})$ itself is rather small, see Figure 5. Around $x=10^{-2}$, the ratio turns positive. At $x=0.3$ it turns negative again, as in the unpolarized case. The complete structure function $g_{1}(x,Q^{2})$ remains positive. The strong negative correction in the small-$x$ region is mainly induced by the sum rule of the vanishing first moment of $\Delta H_{g}^{\rm S}$.
IV Conclusions
@@
In this Letter, we present for the first time numerical results for the single-mass three-loop corrections to the unpolarized structure function $F_{2}(x,Q^{2})$ and the polarized structure function $g_{1}(x,Q^{2})$ at twist-2. The asymptotic analytic results on the heavy-flavor Wilson coefficients are correlated to the kinematic region in which the higher-twist contributions can be neglected. Our numerical results illustrate the impact of the different Wilson coefficients and show the changes in the structure functions going from the $O(a_{s})$ to the $O(a_{s}^{3})$ corrections. In the unpolarized case, the corrections are large in the small-$x$ region, which is also observed in the polarized case. Here, however, the large corrections emerge in a region where the structure function is small, unlike in the former case. The present results are an important step to allow for consistent QCD analyses of the deep-inelastic World data at the level of NNLO, given the partial results used before for combined singlet/non-singlet data analyses. Our new results will allow to improve the knowledge on the strong coupling constant $a_{s}(M_{Z}^{2})$, the PDFs and the charm mass, extracted from the deep-inelastic World data.
@@
Acknowledgements. We would like to thank I. Bierenbaum, A. Goedicke (neé Hasselhuhn), S. Klein, C. Raab, M. Round, M. Saragnese and F. Wißbrock for collaboration in earlier projects and thank P. Marquard for discussions. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP), the UZH Postdoc Grant, grant no. [FK-24-115] and by the Austrian Science Fund (FWF) Grant-DOI 10.55776/P20347.

### v2

@@ 3 The Structure Function $g_{1}(x,Q^{2})$
Figure 6: The different heavy-flavor contributions due to charm quarks to the structure function $g_{1}(x,Q^{2})$ due to the Wilson coefficients $\Delta H_{g}^{\rm S},\Delta H_{q}^{\rm PS},\Delta L_{q}^{\rm NS},\Delta L_{g}^{\rm S}$ and $\Delta L_{g}^{\rm PS}$ at $Q^{2}=100\penalty 10000\ \rm GeV^{2}$ in the strong coupling constant up to $O(a_{s}),O(a_{s}^{2})$ and $O(a_{s}^{3})$.
parton distribution functions, are used in this scheme, $g_{1}(x,Q^{2})$, as a scheme independent quantity, can still be assembled.
For the process of deep-inelastic scattering, the Larin scheme is a consistent scheme, given the number of contributing $\gamma_{5}$ Dirac matrices. A set of Larin-scheme parton distribution functions has been generated in Ref. [61] and is used in the following illustrations. The massless and massive two-loop corrections for the structure function $g_{1}(x,Q^{2})$ were calculated in Ref. [34].
In Figure 5, we illustrate the scale evolution of the complete structure function $g_{1}^{\rm total}(x,Q^{2})$ from $Q^{2}=25\penalty 10000\ \rm GeV^{2}$ to $Q^{2}=1000\penalty 10000\ \rm GeV^{2}$ at three-loop order. With growing values of $Q^{2}$, $xg_{1}^{\rm total}(x,Q^{2})$ grows below $x\sim 0.2$ and depletes at larger values of $x$. Also below $x\sim 10^{-2}$ the order in the scales reverts. The structure functions $g_{1}^{p,d}(x,Q^{2})$ are positive within errors, cf. Ref. [62].
Figure 6 illustrates the contributions of the different heavy-flavor Wilson coefficients to $g_{1}^{c}$ for $Q^{2}=100\penalty 10000\ \rm GeV^{2}$, also with their contributions up to a given order in $a_{s}$. One clearly sees the importance of the higher-order corrections. The largest contribution is due to the gluonic Wilson coefficient $\Delta H_{g}^{\rm S}$. As has been discussed in Refs. [32, 34], the first moment of $\Delta H_{g}^{\rm S}(x,Q^{2})$ vanishes,
$\int_{0}^{1}dx\Delta H_{g}^{\rm S}(x,Q^{2})=0.$ (17)
@@
In Figure 8, we show the ratio Eq. (16) of the charm and bottom contributions to the whole structure function $g_{1}(x,Q^{2})$ at NNLO for $Q^{2}=25,100$ and $1000\penalty 10000\ \rm GeV^{2}$. In the small-$x$ region, the corrections are negative. The largest negative values are about $10\penalty 10000\ \%,-25\penalty 10000\ \%$ and $95\penalty 10000\ \%$ from $Q^{2}=25,100$ to $1000\penalty 10000\ \rm GeV^{2}$. These large corrections occur in a region in which $xg_{1}(x,Q^{2})$ itself is rather small, see Figure 5. Around $x=10^{-2}$, the ratio turns positive. At $x=0.3$ it turns negative again, as in the unpolarized case. The complete structure function $g_{1}(x,Q^{2})$ remains positive. The strong negative correction in the small-$x$ region is mainly induced by the sum rule of the vanishing first moment of $\Delta H_{g}^{\rm S}$.
4 The numerical code
We provide the Fortran-library WILS3HQ of the massive Wilson coefficients to three-loop order as a fast and precise numerical implementation. It is well suited for QCD analyses of deep-inelastic structure functions in the unpolarized and polarized case. The Wilson coefficients are represented as polynomials in the parameters $a_{s},N_{F}$, $\ln(Q^{2}/\mu^{2}),\ln(m^{2}/\mu^{2})$ and flav, see below. Also the $x$-dependence is described by elementary functions and polynomial interpolations. The order in the coupling constant is selected by the parameter IO = 1,2,3. Both for the structure functions $F_{2}$ and $g_{1}$ five different Wilson contribute. In the non-singlet cases $L_{q,2(1)}^{\rm NS}$ there are two distribution valued contributions in addition to the regular part. Likewise, for the Wilson coefficients $H_{2(1),q}^{\rm PS}$ and $L_{2(1),q}^{\rm PS}$, $\delta(1-x)$-terms contribute. The regular terms are represented by suitably deep analytic small- and large-$x$ expansions and grid interpolations for the remainder parts for $x\in[0,1]$, optimized for $x\geq x_{\rm min}=10^{-6}$. The grid interpolations, performed by third order splines [63], are based on 501 points. [^7] We thank S. Kumano and M. Miyama of the AAC-collaboration for allowing us to use their interpolation routines. The provided spline representation can be addressed directly by setting IDAT = 1. The ++-and $\delta$-distribution parts are given in analytic form. The representations need only elementary functions and we used code optimization [64]. The Mellin convolutions with the parton distributions have to be performed in the respective implementations by the users. The data-set initialization is done by the routines INIDAT1 ... 10, where the order of the Wilson coefficients is H2gS, H2qPS, L2qNS, L2gS, L2qPS; HG1gS, HG1qPS, LG1qNS, LG1gS, LG1qPS.
For the grid-files we use the following name convention, illustrated by the example
${\tt L2qPS30011.dat}.$ (18)
The name starts with the name of the Wilson coefficient. The number sequence denotes the order in $a_{s}$, the power of $\ln(m^{2}/\mu^{2})$, the power of $\ln(Q^{2}/\mu^{2})$, the presence of the factor $N_{F}$, the presence of the factor $\tt flav$. The set in Eq. (18) corresponds to the grid for the massive Wilson coefficient $L_{2q}^{\rm PS(3)}$ of the structure function $F_{2}$, proportional to the factor $a_{s}^{3}N_{F}{\tt flav}$.
charm bottom
$L_{q}^{{\rm NS}}$ 22 1-1
$L_{q}^{{\rm PS}}$ $\frac{2}{5}$ $\frac{2}{7}$
$L_{g}^{{\rm S}}$ $\frac{2}{5}$ $\frac{1}{7}$
$H_{q}^{{\rm PS}}$ $\frac{2}{5}$ $\frac{2}{7}$
$H_{g}^{{\rm S}}$ $\frac{2}{5}$ $\frac{1}{7}$
Table 1: Numerical values of the factor flav for the perturbative generation of a heavy charm or bottom quark.
For the charm quark we find the following formulae for the flavor factors [^8] These qunatities are related to the numbers ${\sf w}_{2}$ and ${\sf w}_{3}$ in Refs. [45, 46].
[…]
with $e_{c}$ the electric charge of the charm quark and we already inserted the charges of the three light quark flavors. The corresponding numbers in the case of bottom are obtained by replacing $e_{c}$ with $e_{b}$. The flavor factors for $L_{q}^{\rm PS}$ and $H_{q}^{\rm PS}$ as well as $L_{g}^{\rm S}$ and $H_{g}^{\rm S}$ are the same, since all flavor factors vanish when considering only the three light quark charges. These flavor factors result from the pieces of the three-loop massless Wilson coefficients in the asymptotic massive Wilson coefficients.
The regular parts of the unpolarized Wilson coefficients are denoted by REG...f, as, e.g., REGL2qNS.f in the flavor non-singlet Wilson coefficient for the structure function $F_{2}$ and, correspondingly, with REGLG1qNS.f for the structure function $g_{1}$. The routines of the ++-distribution parts are named by PLU...f, and the coefficients of the $\delta(1-x)$ terms by DEL...f. The argument list is
${\tt(IO,X,as,LM,LQ,NF,flav)},$ (24)
where ${\tt as}=a_{s}(\mu^{2}),{\tt LM}=\ln(m^{2}/\mu^{2}),{\tt LQ}=\ln(Q^{2}/\mu^{2}),{\tt NF}=N_{F},{\tt X}=x$, the Bjorken variable. In the case of the $\delta(1-x)$ terms X is not part of the argument list. For the regular parts there are the auxiliary functions SX....f, LX....f, and GR....f, for the small- and large-$x$ expansions, and the grid part.
The different functions of the library can be compiled with gfortran. We provide ten programs MAIN1.f to MAIN10.f to test the set-up of the respective Wilson coefficients by comparing with numerical results obtained by our initial (semi)-analytic calculation, for some specific choice of parameters.
We add the numerical integration routine DAIND of Ref. [65] to perform the convolution integrals with the parton densities for the convenience of the user. The license conditions to use the library WILS3HQ are to quote the papers Refs. [12, 14, 19, 15, 16, 20, 21, 17, 18] and the present paper, in which the single-mass three-loop corrections were calculated.
5 Conclusions
@@
In this paper, we present for the first time numerical results for the single-mass three-loop heavy-flavor corrections to the unpolarized structure function $F_{2}(x,Q^{2})$ and the polarized structure function $g_{1}(x,Q^{2})$ at twist-2. The asymptotic analytic results on the heavy-flavor Wilson coefficients are correlated to the kinematic region in which the higher-twist contributions can be neglected. Our numerical results illustrate the impact of the different Wilson coefficients and show the changes in the structure functions going from the $O(a_{s})$ to the $O(a_{s}^{3})$ corrections. In the unpolarized case, the corrections are large in the small-$x$ region, which is also observed in the polarized case. Here, however, the large corrections emerge in a region where the structure function is small, unlike in the former case. The present results are an important step to allow for consistent QCD analyses of the deep-inelastic World data at the level of NNLO, given the partial results used before for combined singlet/non-singlet data analyses. Our new results will allow to improve the accuracy of the strong coupling constant $a_{s}(M_{Z}^{2})$, the PDFs and the charm mass, extracted from the deep-inelastic World data.
A fast and numerically precise public Fortran code for the calculation of the asymptotic single-mass heavy-flavor Wilson coefficients, WILS3HQ, for the use in experimental data analyses, is provided as an ancillary file to this paper.
@@
Acknowledgments. We would like to thank I. Bierenbaum, A. Goedicke (né Hasselhuhn), S. Klein, C. Raab, M. Round, M. Saragnese and F. Wißbrock for collaboration in earlier projects and thank P. Marquard for discussions. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP), the UZH Postdoc Grant, grant no. [FK-24-115] and by the Austrian Science Fund (FWF) Grant-DOI 10.55776/P20347.